OpenStudy (anonymous):

Write answer in Decimal form 3/4 % . 60

OpenStudy (anonymous):

It means: \[\frac{3}{4} \% \; of \; 60 \implies \frac{\frac{3}{4}}{100} \times 60 \implies \frac{3}{400} \times 60 = ??\]